# Probing the equilibration of the QCD matter created in heavy-ion collisions with dileptons

### v1

@@ I Introduction
Creating and measuring the properties of Quark-Gluon Plasma (QGP), an exotic state of nuclear matter composed of colored quarks and gluons obeying Quantum Chromodynamics (QCD), is mainstream research in subatomic physics. This phase existed only microseconds after the Big Bang Jacak and Muller (2012), could well constitute the inner core of massive stellar objects Annala et al. (2023), and occupies a good fraction of the QCD phase diagram. To simulate the dense and high-temperature environment of the early universe in laboratory conditions, a practical approach is to collide two nuclei with extremely high collision energies, such as what is currently being performed at the Relativistic Heavy Ion Collider (RHIC) and the Large Hadron Collider (LHC). An important milestone in the understanding and modeling of the QGP is the realization that the QGP fireball behaves like a fluid Gale et al. (2013) characterized by an extremely small specific shear viscosity Policastro et al. (2001). This is measured through correlations of final-state particles, more specifically by examining anisotropies in the azimuthal momentum distribution of the final charged hadrons Poskanzer and Voloshin (1998). Therefore, the charged hadrons production and hadronic flow observables do carry information about the QGP medium. This degree of understanding has been achieved through years of developing a variety of theoretical approaches, evolving into what is currently known as multistage modeling.
@@
Whereas the information carried by hadrons is dominated by the conditions which exist at the freeze-out hypersurface, electromagnetic (EM) radiation is emitted throughout the space-time evolution and exits the hadronic medium. This is true for photons Gale et al. (2022); Paquet et al. (2016); van Hees et al. (2011) and for lepton pairs Gale and Haglin (2003); Gale and Kapusta (1987); van Hees and Rapp (2008). This class of observables is therefore complementary to that constituted by strongly interacting particles. The EM probes carry information about the their space-time point of production, which is typically earlier than for hadronic production. In view of those attributes, to calculate the yield of electromagnetic radiation it is essential to accurately simulate the dynamical evolution history and to account for the different stages in a relativistic heavy ion collision, using a comprehensive hybrid framework. Microscopically, it is also crucial to precisely calculate the emission rate of EM probes. For instance, recent studies have considered modifying the emission rates of EM probes to include viscosity corrections Vujanovic et al. (2020, 2016, 2018, 2014), corrections owing to the anisotropy of the emitting medium Hauksson et al. (2018); Kasmaei and Strickland (2019) or next-to-leading-order at finite baryonic density Churchill et al. (2024a).
@@
From the perspective of simulating the space-time evolution of the medium, the currently accepted minimum standard of multistage modeling Schenke et al. (2010, 2011); Paquet et al. (2016); Pang et al. (2018); Nijs et al. (2021); Wu et al. (2022); Du et al. (2024) includes fluctuating initial conditions, relativistic hydrodynamic evolution, and a hadronic afterburner. Models following this prescription have been used to predict and interpret a wide variety of hadronic observables. Considering the different stages involved in this multistage model construction, the question of how “hydrodynamization” can be achieved in the very short time demanded by the interpretation of data ( fm/) still remains somewhat mysterious, even though progress in the understanding of early time dynamics has been made in the last few years. For example, some effort has been devoted to studying the dynamical evolution of the pre-equilibrium (pre-hydrodynamics) stage, going from an initial state (either Glauber+free-streaming, or IP-Glasma Schenke et al. (2012a, b)) to viscous fluid dynamics. Alternatively, solving effective kinetic theories with anisotropic initial gluon distributions offers another approach to simulate the out-of-equilibrium stage Arnold et al. (2003); Du and Schlichting (2021); Garcia-Montero et al. (2024a), at the price of solving a set of coupled relativistic Boltzmann equations. However, recent work utilizing linear response theory offers a promising alternative to the brute force solution of the full non-linear Boltzmann equation, and propagates the energy-momentum tensor forward in time to be finally used as an initial state to the fluid-dynamical stage Kurkela et al. (2019a, b). That model – KøMPøST – will be used later in this work to simulate the approach to kinetic equilibrium.
@@
This paper is organized as follows: The next section briefly introduces the multistage model and details of the calculations of dilepton production and dilepton anisotropic flow from NLO emission rates. In Section III, we first validate the model calibration using charged hadrons observables. We then explore in detail the effects of the pre-equilibrium stage and chemical equilibrium on dilepton production and dilepton anisotropic flow. A summary and outlook are provided in Section IV.
@@ II.1 Multistage modeling
In this work, the multistage modeling of heavy ion collisions is accomplished by the iEBE-MUSIC hydrodynamic approach Schenke et al. (2010, 2011); Paquet et al. (2016), supplemented by IP-Glasma initial states Schenke et al. (2012a, b), iSS particlization, and a UrQMD hadron afterburner Shen et al. (2016); Bass et al. (1998). Since one of the goals of this study is to examine the contribution of the early-time dynamics, the hybrid model also incorporates a non-equilibrium approach to the pre-hydrodynamic phase: KøMPøST .
@@
Starting from the initialization, the primordial evolution of the system relies on IP-Glasma (the Impact-Parameter-dependent Glasma), which samples nucleon positions from the Woods-Saxon distribution, with each nucleon modeled as three constituent quarks. Upon establishing the initial configuration of these quarks, the saturation scale is determined via the IP-SAT model (impact parameter dependent dipole saturation model) Kowalski and Teaney (2003), which is calibrated using data from deeply inelastic scattering experiments at HERA Rezaeian et al. (2013). This saturation scale correlates with the variance in the color charge density distribution. These serve as the sources for the small- classical gluon field within the Color Glass Condensate (CGC) framework. The evolution of the gluon field is then described by the classical Yang-Mills (CYM) equations. Following the collision of two gluon fields, the field continues to evolve according to the CYM equations until the initial proper time of the pre-equilibrium evolution111EKT stands for “Effective Kinetic Theory”, the dynamical picture underlying KøMPøST .. In this work, the initial proper time is taken to be 0.1 fm/ Kurkela et al. (2019b); Gale et al. (2022), which corresponds to a saturation scale GeV.
@@
At the initial proper time of the pre-equilibrium phase, the complete energy-momentum tensor constructed from IP-Glasma constitutes the input of KøMPøST , as done in Ref. Gale et al. (2022). For detailed discussions of KøMPøST , refer to Kurkela et al. (2019b, a). In this study, the end of the pre-equilibrium phase or initial hydrodynamics time will be = 0.8 fm/, to be consistent with previous estimates Kurkela et al. (2019b); Gale et al. (2022). In the future, the proper time interval for pre-equilibrium could be determined through Bayesian inference Everett et al. (2021a); Heffernan et al. (2023).
@@
Following the pre-equilibrium evolution, the strongly interacting system evolves and is governed by relativistic hydrodynamics. In MUSIC, the initial energy density and initial flow velocity are initialized through Landau matching condition , and the complete initial energy-momentum tensor is that passed on from KøMPøST . The dissipative parts of the energy-momentum tensor – the shear tensor and bulk pressure – are initialized via
@@
where and denote the ideal part of the energy-momentum tensor and pressure in the hydrodynamics phase, respectively222KøMPøST is conformal in its current formulation.. Combined with the HotQCD lattice equation of state (EOS) Bazavov et al. (2014), the energy-momentum conservation and the second-order Israel-Stewart relaxation-type equations are solved numerically Schenke et al. (2010). In this study, the specific shear viscosity is taken as 0.12, as it is in the KøMPøST evolution. The bulk viscosity is modeled as temperature-dependent with a double Gaussian parameterization; for the details, please consult Ref. Gale et al. (2022). The hydrodynamics evolution continues until the local energy density drops to = 0.18 GeV/fm3. Subsequently, the Cooper-Frye prescription Cooper and Frye (1974) is used to convert the macroscopic thermodynamic variables of the hydrodynamics phase into the phase-space information of microscopic particles, via the iSS sampler module. During the sampling process, corrections for shear and bulk viscous components are applied according to the 14-moment method Ryu et al. (2015); Schenke et al. (2020). Finally, the thermal hadrons are fed into the UrQMD model Shen et al. (2016) to continue through hadronic cascading, including decay and formation channels.
@@ II.2 Dilepton production
The differential thermal emission rate of lepton pairs of energy , three-momentum , and invariant mass is written as [See; forexample; ][; andreferencestherein.]Kapusta:2023eix
@@
Once the dilepton emission rate is determined, the dilepton production can be evaluated by integrating over the space-time volume which is provided by the dynamical evolution of the QGP phase or pre-equilibrium phase. The differential dilepton spectrum is
As written, the differential emission rate is evaluated in the LRF of the fluid cell. However, dileptons are measured in the lab frame, which is denoted with a prime in Eq. (4) and (5). Therefore, when performing the above integration, the dilepton momentum in the lab frame, , must be boosted into the LRF of each flow cell using , where is the Lorentz matrix which depends on the local flow velocity . For the sake of simplicity, the prime notation, which indicates the measured dilepton observables in the lab frame, and boost notation are omitted for the remainder of this paper.
Proceeding beyond the dilepton spectra, the -th order dilepton flow vector can also be calculated from the fully differential emission rate in a single event, as defined in Paquet et al. (2016); Vujanovic et al. (2016, 2018), specifically
@@ III Results
We shall concentrate on the analysis of dilepton production and anisotropic flow in Pb+Pb collisions at a collision energy of TeV. After our model is calibrated using hadronic flow observables, we will perform an analysis of the effects of chemical equilibration on dilepton observables. The dependence on both transverse momentum and invariant mass of dilepton will be considered.
@@ III.1 Hadrons
In Fig. 1(a), the hadron multiplicity as a function of centrality bins in the central rapidity region for TeV Pb+Pb collisions is shown. The experimental data is from the ALICE collaboration Acharya et al. (2020); Adam et al. (2016). The multiplicity is used to set the overall normalization of the initial energy-momentum tensor. It can be seen that the centrality dependence of the data for both charged particles and for identified hadrons is well reproduced. Note that the centrality selection in our study is determined from the final charged hadrons within the pseudorapidity range . This selection method is similar to what is used in the experimental analysis, allowing for an appropriate comparison between the model results and data.
@@
In Fig. 1(b), the mean transverse momentum of identified hadrons is shown as a function of centrality bins, together with that of all charged hadrons. It can be observed that our results agree well with ALICE data for mesons (and also for all charged particles) up to the 75% centrality bin. For protons produced in central collisions, the average transverse momentum is slightly underestimated () as was found in a recent similar study Schenke et al. (2020). As was concluded there, future work will address this discrepancy; it is known that protons are especially sensitive to the details of the bulk viscosity Ryu et al. (2015). For peripheral centralities, results and data deviate, with the model rising about the data as the centrality grows. This is understood as follows: our current model includes both pre-equilibrium (KøMPøST ) and hydrodynamic (MUSIC) evolutions, assuming a constant lifetime for the pre-equilibrium stage ( 0.7 fm/) across all centrality bins. In central collisions, the lifetime of hydrodynamic stage is significantly longer ( fm/) than that of the pre-equilibrium stage. However, in peripheral collisions, the lifetime of the hydro stage becomes smaller. As KøMPøST and IP-Glasma both assume that the collision system has conformal symmetry Kurkela et al. (2019b) and thus vanishing bulk viscosity, the hydro phase has a comparatively shorter time to slow down the transverse expansion as compared with the situation in central collisions. In future studies, the lifetime of the pre-equilibrium stage could be considered as dependent on the initial energy density and centrality class, and could be determined via Bayesian inference based on experimental data Everett et al. (2021b).
@@
In Fig. 1(c), the centrality dependence of different order anisotropic flows for final charged hadrons is presented. The flow coefficients are calculated using the two-particle correlation method, with pseudorapidity and transverse momentum GeV. Calculation results describe the experimental data up to the intermediate centrality bins but begin to deviate from the data beyond 45% centrality, due to the relatively large contribution and deficiency of bulk viscosity in the early-time dynamics. Finally combined with the multiplicity, mean transverse momentum , and flow coefficient of hadrons, it can be concluded that different hadronic observables present varying sensitivities to pre-equilibrium stages, but that the global representation of hadronic observables is more than adequate for our purposes.
@@ III.2 Dileptons
We estimate the dilepton production in the pre-equilibrium stage by using the same emission rate as in the hydrodynamic stage. The local “temperature” and fluid velocity can be obtained through the Landau matching method for the energy-momentum tensor of pre-equilibrium stages with the HotQCD EOS. Thus, using the equilibrium emission rate is but an effective method to estimate dilepton production of the pre-equilibrium stage. In addition, using the same EOS has the advantage of ensuring a smooth connection between the pre-equilibrium and hydrodynamic stagesGale et al. (2022). Note that the combination of (2+1)-D hydrodynamics with the pre-equilibrium simulation is utilized to estimate the evolution of medium background at spatial rapidity =0. However, dileptons may be emitted from regions with finite spatial rapidity Paquet and Bass (2022). Therefore, in Eq. (4) and Eq. (5), spatial rapidity integration is performed by assuming the local temperature and local flow velocity to be independent of spatial rapidity within a narrow window, selecting in this work to evaluate observables at .
@@
In the IP-Glasma model and KøMPøST model, both models assume that the collision system is purely gluonic. However, dilepton production originates from quark-antiquark annihilation at LO or, e.g., Compton scattering at NLO. This demands the presence of quarks and antiquarks in the collision system. Naturally, starting from all gluons, quarks and antiquarks will gradually be produced via process, leading to the establishment of chemical equilibrium. Previous studies of chemical equilibration in heavy ion collisions have been performed with EKT, and have defined a chemical relaxation time, Kurkela and Mazeliauskas (2019a, b). We adopt this definition here and model an effective suppression factor (SF):
@@
This suppression factor multiplies the dilepton emission rate and represents the fact that the fermions are under-populated as compared to the situation in chemical equilibrium. Since our rates are derived up to, and including NLO, we allow for flexibility by studying cases where the emission are either linear and quadratic in the SF, in addition to the case where SF=1. Inspired by the work in Refs. Kurkela and Mazeliauskas (2019b, a), where the time for chemical equilibrium is found to be , the constant in Eq. (9) is determined by requiring , as in those references. Note that the suppression factor thusly generated resembles that in Ref. Gale et al. (2022) – a similar study as that performed here, but for photons – with the difference that the approach to equilibrium is temperature-dependent here.
@@
In Fig. 2, the suppression factor is presented as a function of proper time for three typical local temperatures: = 0.1, 0.2, and 0.3 GeV, corresponding to chemical equilibrium times of 1.19 fm, 1.79 fm, and 3.57 fm, respectively. The shear viscosity per unit entropy density is set at 0.12, consistent with our simulation of the pre-equilibrium and hydrodynamic stages. From the Fig. 2, it is evident that systems with higher temperatures reach chemical equilibrium faster than those at lower temperatures. The time of chemical equilibrium may be slightly longer than the initial proper time of hydrodynamic evolution Kurkela and Mazeliauskas (2019a). We will now examine the details of the production of lepton pair using our multistage model,, with a dynamical pre-equilibrium stage.

### v2

@@ I Introduction
Creating and measuring the properties of quark-gluon plasma (QGP), an exotic state of nuclear matter composed of colored quarks and gluons obeying quantum chromodynamics (QCD), is mainstream research in subatomic physics. This phase existed only microseconds after the big bang Jacak and Muller (2012), could well constitute the inner core of massive stellar objects Annala et al. (2023), and occupies a good fraction of the QCD phase diagram. To simulate the dense and high-temperature environment of the early universe in laboratory conditions, a practical approach is to collide two nuclei with extremely high collision energies, as is currently being performed at the Relativistic Heavy Ion Collider (RHIC) and the Large Hadron Collider (LHC). An important milestone in the understanding and modeling of the QGP is the realization that the QGP fireball behaves like a fluid Gale et al. (2013) characterized by an extremely small specific shear viscosity Policastro et al. (2001). This is measured through correlations of final-state particles, more specifically by examining anisotropies in the azimuthal momentum distribution of the final charged hadrons Poskanzer and Voloshin (1998). Therefore, the charged hadrons production and hadronic flow observables do carry information about the QGP medium. This degree of understanding has been achieved through years of developing a variety of theoretical approaches, evolving into what is currently known as multistage modeling.
@@
Whereas the information carried by hadrons is dominated by the conditions which exist at the freeze-out hypersurface, electromagnetic (EM) radiation is emitted throughout the space-time evolution and exits the hadronic medium. This is true for photons Gale et al. (2022); Paquet et al. (2016); van Hees et al. (2011) and for lepton pairs Gale and Haglin (2003); Gale and Kapusta (1987); van Hees and Rapp (2008). This class of observables is therefore complementary to that constituted by strongly interacting particles. The EM probes carry information about the their space-time point of production, which is typically earlier than for hadronic production. In view of those attributes, to calculate the yield of electromagnetic radiation it is essential to accurately simulate the dynamical evolution history and to account for the different stages in a relativistic heavy ion collision, using a comprehensive hybrid framework. Microscopically, it is also crucial to precisely calculate the emission rate of EM probes. For instance, recent studies have considered modifying the emission rates of EM probes to include viscosity corrections Vujanovic et al. (2020, 2016, 2018, 2014), corrections owing to the anisotropy of the emitting medium Hauksson et al. (2018); Kasmaei and Strickland (2019) or next-to-leading order (NLO) results at finite baryonic density Churchill et al. (2024a).
@@
From the perspective of simulating the space-time evolution of the medium, the currently accepted minimum standard of multistage modeling Schenke et al. (2010, 2011); Paquet et al. (2016); Pang et al. (2018); Nijs et al. (2021); Wu et al. (2022); Du et al. (2024) includes fluctuating initial conditions, relativistic hydrodynamic evolution, and a hadronic afterburner. Models following this prescription have been used to predict and interpret a wide variety of hadronic observables. Considering the different stages involved in this multistage model construction, the question of how “hydrodynamization” can be achieved in the very short time demanded by the interpretation of data ( fm/) still remains somewhat mysterious, even though progress in the understanding of early time dynamics has been made in the last few years. For example, some effort has been devoted to studying the dynamical evolution of the preequilibrium (pre-hydrodynamics) stage, going from an initial state (either Glauber+free-streaming, or IP-Glasma Schenke et al. (2012a, b)) to viscous fluid dynamics. Alternatively, solving effective kinetic theories with anisotropic initial gluon distributions offers another approach to simulate the out-of-equilibrium stage Arnold et al. (2003); Du and Schlichting (2021); Garcia-Montero et al. (2024a), at the price of solving a set of coupled relativistic Boltzmann equations. However, recent work utilizing linear response theory offers a promising alternative to the brute force solution of the full non-linear Boltzmann equation, and propagates the energy-momentum tensor forward in time to be finally used as an initial state to the fluid-dynamical stage Kurkela et al. (2019a, b). That model – KøMPøST – will be used later in this work to simulate the approach to kinetic equilibrium.
@@
This paper is organized as follows: The next section briefly introduces the multistage model and details of the calculations of dilepton production and dilepton anisotropic flow from NLO emission rates. In Sec. III, we first validate the model calibration using charged hadrons observables. We then explore in detail the effects of the preequilibrium stage and chemical equilibrium on dilepton production and dilepton anisotropic flow. A summary and outlook are provided in Sec. IV.
@@ II.1 Multistage modeling
In this work, the multistage modeling of heavy ion collisions is accomplished by the iEBE-MUSIC hydrodynamic approach Schenke et al. (2010, 2011); Paquet et al. (2016), supplemented by IP-Glasma initial states Schenke et al. (2012a, b), iSS particlization, and a UrQMD hadron afterburner Shen et al. (2016); Bass et al. (1998). Since one of the goals of this study is to examine the contribution of the early-time dynamics, the hybrid model also incorporates a nonequilibrium approach to the prehydrodynamic phase: KøMPøST .
@@
Starting from the initialization, the primordial evolution of the system relies on IP-Glasma (the impact-parameter-dependent glasma), which samples nucleon positions from the Woods-Saxon distribution, with each nucleon modeled as three constituent quarks. Upon establishing the initial configuration of these quarks, the saturation scale is determined via the IP-SAT model (impact-parameter-dependent dipole saturation model) Kowalski and Teaney (2003), which is calibrated using data from deeply inelastic scattering experiments at HERA Rezaeian et al. (2013). This saturation scale correlates with the variance in the color charge density distribution. These serve as the sources for the small- classical gluon field within the color glass condensate (CGC) framework. The evolution of the gluon field is then described by the classical Yang-Mills (CYM) equations. Following the collision of two gluon fields, the field continues to evolve according to the CYM equations until the initial proper time of the preequilibrium evolution111EKT stands for “effective kinetic theory,” the dynamical picture underlying KøMPøST .. In this work, the initial proper time is taken to be 0.1 fm/ Kurkela et al. (2019b); Gale et al. (2022), which corresponds to a saturation scale GeV.
@@
At the initial proper time of the preequilibrium phase, the complete energy-momentum tensor constructed from IP-Glasma constitutes the input of KøMPøST , as done in Ref. Gale et al. (2022). For detailed discussions of KøMPøST , refer to Kurkela et al. (2019b, a). In this study, the end of the preequilibrium phase or initial hydrodynamics time will be = 0.8 fm/, to be consistent with previous estimates Kurkela et al. (2019b); Gale et al. (2022). In the future, the proper time interval for preequilibrium could be determined through Bayesian inference Everett et al. (2021a); Heffernan et al. (2023).
@@
Following the preequilibrium evolution, the strongly interacting system evolves and is governed by relativistic hydrodynamics. In MUSIC, the initial energy density and initial flow velocity are initialized through Landau matching condition , and the complete initial energy-momentum tensor is that passed on from KøMPøST . The dissipative parts of the energy-momentum tensor – the shear tensor and bulk pressure – are initialized via
@@
where and denote the ideal part of the energy-momentum tensor and pressure in the hydrodynamics phase, respectively222KøMPøST is conformal in its current formulation.. Combined with the HotQCD Collaboration’s lattice equation of state (EOS) Bazavov et al. (2014), the energy-momentum conservation and the second-order Israel-Stewart relaxation-type equations are solved numerically Schenke et al. (2010). In this study, the specific shear viscosity is taken as 0.12, as it is in the KøMPøST evolution. The bulk viscosity is modeled as temperature-dependent with a double Gaussian parametrization; for the details, please consult Ref. Gale et al. (2022). The hydrodynamics evolution continues until the local energy density drops to = 0.18 GeV/fm3. Subsequently, the Cooper-Frye prescription Cooper and Frye (1974) is used to convert the macroscopic thermodynamic variables of the hydrodynamics phase into the phase-space information of microscopic particles, via the iSS sampler module. During the sampling process, corrections for shear and bulk viscous components are applied according to the 14-moment method Ryu et al. (2015); Schenke et al. (2020). Finally, the thermal hadrons are fed into the UrQMD model Shen et al. (2016) to continue through hadronic cascading, including decay and formation channels.
@@ II.2 Dilepton production
Considering three light flavors (), the differential thermal emission rate of lepton pairs of energy , three-momentum , and invariant mass is written as [See; forexample; ][; andreferencestherein.]Kapusta:2023eix
@@
Once the dilepton emission rate is determined, the dilepton production can be evaluated by integrating over the space-time volume which is provided by the dynamical evolution of the QGP phase or preequilibrium phase. The differential dilepton spectrum is
As written, the differential emission rate is evaluated in the LRF of the fluid cell. However, dileptons are measured in the laboratory frame, which is denoted with a prime in Eqs. (4) and (5). Therefore, when performing the above integration, the dilepton momentum in the laboratory frame, , must be boosted into the LRF of each flow cell using , where is the Lorentz matrix which depends on the local flow velocity . For the sake of simplicity, the prime notation, which indicates the measured dilepton observables in the laboratory frame, and boost notation are omitted for the remainder of this paper.
Proceeding beyond the dilepton mass spectra, the th order dilepton flow vector can also be calculated from the fully differential emission rate in a single event, as defined in Paquet et al. (2016); Vujanovic et al. (2016, 2018), specifically
@@ III Results
We shall concentrate on the analysis of dilepton production and anisotropic flow in Pb+Pb collisions at a collision energy of TeV. After our model is calibrated using hadronic flow observables, we will perform an analysis of the effects of chemical equilibration on dilepton observables. The dependence on both transverse momentum and invariant mass of dileptons will be considered.
@@ III.1 Hadrons
In Fig. 1(a), the hadron multiplicity as a function of centrality bins in the central rapidity region for TeV Pb+Pb collisions is shown. The experimental data are from the ALICE Collaboration Acharya et al. (2020); Adam et al. (2016). The multiplicity is used to set the overall normalization of the initial energy-momentum tensor. It can be seen that the centrality dependence of the data for both charged particles and for identified hadrons is well reproduced. Note that the centrality selection in our study is determined from the final charged hadrons within the pseudorapidity range . This selection method is similar to what is used in the experimental analysis, allowing for an appropriate comparison between the model results and data.
@@
In Fig. 1(b), the mean transverse momentum of identified hadrons is shown as a function of centrality bins, together with that of all charged hadrons. It can be observed that our results agree well with ALICE data for mesons (and also for all charged particles) up to the 75% centrality bin. For protons produced in central collisions, the average transverse momentum is slightly underestimated () as was found in a recent similar study Schenke et al. (2020). As was concluded there, future work will address this discrepancy; it is known that protons are especially sensitive to the details of the bulk viscosity Ryu et al. (2015). For peripheral centralities, results and data deviate, with the model rising above the data as the centrality grows. This is understood as follows: our current model includes both preequilibrium (KøMPøST ) and hydrodynamic (MUSIC) evolutions, assuming a constant lifetime for the preequilibrium stage ( 0.7 fm/) across all centrality bins. In central collisions, the lifetime of the hydrodynamic stage is significantly longer [ fm/] than that of the preequilibrium stage. However, in peripheral collisions, the lifetime of the hydro stage becomes smaller. As KøMPøST and IP-Glasma both assume that the collision system has conformal symmetry Kurkela et al. (2019b) and thus vanishing bulk viscosity, the hydro phase has a comparatively shorter time to slow down the transverse expansion as compared with the situation in central collisions. In future studies, the lifetime of the preequilibrium stage could be considered as dependent on the initial energy density and centrality class, and could be determined via Bayesian inference based on experimental data Everett et al. (2021b).
@@
In Fig. 1(c), the centrality dependence of different order anisotropic flows for final charged hadrons is presented. The flow coefficients are calculated using the two-particle correlation method, with pseudorapidity and transverse momentum GeV. Calculation results describe the experimental data up to the intermediate centrality bins but begin to deviate from the data beyond 45% centrality, due to the relatively large contribution and deficiency of bulk viscosity in the early-time dynamics. Finally combined with the multiplicity, mean transverse momentum , and flow coefficient of hadrons, it can be concluded that different hadronic observables present varying sensitivities to preequilibrium stages, but that the global representation of hadronic observables is more than adequate for our purposes.
@@ III.2 Dileptons
We estimate the dilepton production in the preequilibrium stage by using the same emission rate as in the hydrodynamic stage. The local “temperature” and fluid velocity can be obtained through the Landau matching method for the energy-momentum tensor of the preequilibrium stage with the HotQCD EOS. Thus, using the equilibrium emission rate is an effective method to estimate dilepton production of the preequilibrium stage. In addition, using the same EOS has the advantage of ensuring a smooth connection between the preequilibrium and hydrodynamic stages Gale et al. (2022). Note that the combination of (2+1)-dimensional hydrodynamics with the preequilibrium simulation is utilized to estimate the evolution of medium background at spatial rapidity =0. However, dileptons may be emitted from regions with finite spatial rapidity Paquet and Bass (2022). Therefore, in Eqs. (4) and (5), spatial rapidity integration is performed by assuming the local temperature and local flow velocity to be independent of spatial rapidity within a narrow window, selecting in this work to evaluate observables at .
@@
In the IP-Glasma model and KøMPøST model, both models assume that the collision system is purely gluonic. However, dilepton production originates from quark-antiquark annihilation at LO or, e.g., Compton scattering at NLO. This demands the presence of quarks and antiquarks in the collision system. Naturally, starting from all gluons, quarks and antiquarks will gradually be produced via the process, leading to the establishment of chemical equilibrium. Previous studies of chemical equilibration in heavy ion collisions have been performed with EKT, and have defined a chemical relaxation time, Kurkela and Mazeliauskas (2019a, b). We adopt this definition here and model an effective suppression factor (SF):
@@
Inspired by the work in Refs. Kurkela and Mazeliauskas (2019b, a), where the time for chemical equilibrium is found to be , the constant in Eq. (9) is determined by requiring , as in those references. This suppression factor multiplies the dilepton emission rate and represents the fact that the fermions are under-populated as compared to the situation in chemical equilibrium. Since our rates are derived up to, and including NLO, we allow for flexibility by studying cases where the emission is either linear or quadratic in the SF, in addition to the case where SF=1. Note that the suppression factor thusly generated resembles that in Ref. Gale et al. (2022) – a study similar to that performed here, but for photons – with the difference that the approach to equilibrium is temperature dependent here.
Obtaining the fermionic content in the non-equilibrium environment from first principles would be eminently desirable. However, while this is possible in principle, it would entail developing a multistage model where a non-equilibrium transport dovetails to a viscous hydro, and where all parameters would be tuned (perhaps through a Bayesian study) to observables; this is the topic of future work. Instead, we use here this flexible suppression factor to explore the sensitivity of the dilepton signal to its specific form.
@@
In Fig. 2, the suppression factor is presented as a function of proper time for three typical local temperatures: = 0.1, 0.2, and 0.3 GeV, corresponding to chemical equilibrium times of 1.19 , 1.79 , and 3.57 fm, respectively. The shear viscosity per unit entropy density is set at 0.12, consistent with our simulation of the preequilibrium and hydrodynamic stages. From the Fig. 2, it is evident that systems with higher temperatures reach chemical equilibrium faster than those at lower temperatures. The time of chemical equilibrium may be slightly longer than the initial proper time of hydrodynamic evolution Kurkela and Mazeliauskas (2019a). We will now examine the details of the production of lepton pair using our multistage model, with a dynamical preequilibrium stage.